what is the general form of 0,2,6,12,20,30,42,.....
Answer by Theo(13342)

this sequence is determined by multiplying the number of the term you're on by the number of the term preceding it.

the formula would state that the value in term n is equal to n * (n-1).

in algebra language that would be An = n * (n-1)

An is the value in term n.
n is the term number.

using this formula, you would get:

A1 = 1 * 0 = 0
A2 = 2 * 1 = 2
A3 = 3 * 2 = 6
A4 = 4 * 3 = 12
A5 = 5 * 4 = 20
A6 = 6 * 5 = 30
etc...